# A four-operator splitting algorithm for nonconvex and nonsmooth optimization

### v1

@@ 3.3 Subsequential convergence
After obtaining the stepsize upper bounds, our next goal is to show subsequential convergence and convergence rates of Algorithm 1. We first establish the boundedness of its iterate sequence generated by Algorithm 1.
@@
    $\tau\geq 2$ satisfies Eqs. 3.47 and 3.48, and $\alpha\in\left(\frac{\tau\mu_{*}}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})},\frac{\tau\mu^{*}}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}\right)\cap\left(0,\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right)$.
@@
By our choice of $(\alpha,\beta,\gamma)$, it follows from Theorems 3.5, 3.9 and 3.12 that $\{V_{k}\}$ is strictly decreasing. Meanwhile, from Remark 3.6 and Remark 3.11 that under conditions (a) and (b), respectively, we have $\alpha<\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$. In condition (c), it is explicitly assumed that $\alpha<\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$. Hence, the second term in (3.54) is always bounded below by zero. By our choice of $\beta$, the last term is also nonnegative. With these, we conclude that $\{\Psi(y^{k})\}$ is bounded above, and therefore $\{y^{k}\}$ is a bounded sequence by the level-boundedness of $\Psi$. On the other hand, since $\Psi$ is closed, $\Psi$ is also bounded below. It thus follows from Eq. 3.54 that $\{{\left\|{y^{k}-x^{k-1}}\right\|}\}$ is bounded above. Since $\{y^{k}\}$ is bounded, it then follows that $\{x^{k-1}\}$ is also bounded. Finally, since $z^{k}=x^{k}+\alpha\nabla f(x^{k})$ by (3.1) and $\nabla f$ is continuous, we obtain the boundedness of $\{z^{k}\}$. This completes the proof of part (i).
@@
$0\in\nabla f(w^{*})+\partial g(w^{*})+\nabla h(w^{*})+\partial p(w^{*}).$
@@
$T_{\Lambda}(y,z)\coloneqq\left\{\begin{pmatrix}y^{\scriptscriptstyle+}\\ z+\tau(y^{\scriptscriptstyle+}-{\rm prox}_{\alpha f}(z))\end{pmatrix}:y^{\scriptscriptstyle+}\in P_{\Lambda}(y,z)\right\}.$
@@
If $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$, then $y^{*}\in P_{\Lambda}(y^{*},z^{*})$ and $z^{*}\in z^{*}+\tau(y^{*}-{\rm prox}_{\alpha f}(z^{*}))$. From the last equation and noting that ${\rm prox}_{\alpha f}$ is single-valued since $\alpha<L_{{}_{\rm{f}}}^{-1}$, we have
@@
Adding (3.57) and (3.58), we get $-\nabla h(y^{*})-\xi^{*}\in\nabla f(y^{*})+\partial g(y^{*})$, proving the desired result. ∎
@@
$\omega\coloneqq\begin{cases}\frac{2\beta(V_{0}-\Psi^{*})}{\min\left\{2\beta\zeta,1-\beta L_{{}_{\rm{p}}}\right\}}&\text{if}~{}\beta<L_{{}_{\rm{p}}}^{-1},\\ \frac{4\gamma(V_{0}-\Psi^{*})}{\min\left\{2\gamma\zeta,1-\gamma\rho_{{}_{\rm{g}}}\right\}}&\text{if}~{}\gamma<\rho_{{}_{\rm{g}}}^{-1},\end{cases}\quad\text{where}\quad\zeta\coloneqq\frac{-c(\alpha)}{2\tau\alpha(1+\alpha L_{{}_{\rm{f}}})^{2}}>0$
@@
Let $\{(x^{k_{j}},y^{k_{j}},z^{k_{j}})\}$ be a subsequence such that $(x^{k_{j}},y^{k_{j}},z^{k_{j}})\to(x^{*},y^{*},z^{*})$. From the proof of Proposition 3.15, we have that ${\left\|{y^{k}-x^{k-1}}\right\|}\to 0$ and ${\left\|{y^{k}-y^{k-1}}\right\|}\to 0$. It then follows that $y^{k_{j}}\to x^{*}$ by the triangle inequality, and therefore $x^{*}=y^{*}$. To show that $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$, we first note that since $x^{*}=y^{*}$ and $z^{k}=x^{k}+\alpha\nabla f(x^{k})$ by (2.4), it follows that $z^{*}=y^{*}+\alpha\nabla f(y^{*})$. Since $\alpha<\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\leq\frac{1}{L_{{}_{\rm{f}}}}$, then $f+\frac{1}{2\alpha}\|\cdot\|^{2}$ is strongly convex by Eq. 2.5 and therefore $y^{*}\in{\rm prox}_{\alpha f}(z^{*})$. Thus, $z^{*}=z^{*}+\tau(y^{*}-{\rm prox}_{\alpha f}(z^{*}))$. Hence, to show that $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$, it remains to prove that $y^{*}\in P_{\Lambda}(y^{*},z^{*})$. To this end, note that by the convexity and continuity of $\frac{L_{{}_{\rm{p}}}}{2}\|\cdot\|^{2}-p$ on $\operatorname*{dom}(g)$ (by Assumption 1.1(d)) and the formula (3.21), we have from (Beck, 2017, Theorem 3.16) and (3.21) that $\{\xi^{k_{j}}\}$ is bounded. Using again the continuity of $p$ on $\operatorname*{dom}(g)$ together with (2.2) and the fact that $\{y^{k}\}\subseteq\operatorname*{dom}(g)$, we see that accumulation points of $\{\xi^{k_{j}}\}$ belong to $\partial p(y^{*})$. Without loss of generality, we may assume that $\xi^{k_{j}}\to\xi^{*}$ where $\xi^{*}\in\partial p(y^{*})$.
@@
Note that $y^{k_{j}+1}\to y^{*}$ and $u^{k_{j}}\to u^{*}\coloneqq\frac{\gamma}{\alpha}(2y^{*}-z^{*}-\alpha\nabla h(y^{*}))+\frac{\gamma}{\beta}(y^{*}-\beta\xi^{*})$, where we have made use of the fact that $\xi^{k_{j}}\to\xi^{*}$ and $x^{*}=y^{*}$. To prove the claim that $y^{*}\in P_{\Lambda}(y^{*},z^{*})$, we only need to show that $y^{*}\in{\rm prox}_{\gamma g}(u^{*})$, noting that $y^{*}\in{\rm prox}_{\alpha f}(z^{*})$
@@
$\leq(f+h)(\bar{x})+\left\langle\nabla(f+h)(\bar{x}),x-\bar{x}\right\rangle+p(\bar{x})+\left\langle\bar{\xi},x-\bar{x}\right\rangle+\frac{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+L_{{}_{\rm{p}}}}{2}{\left\|{x-\bar{x}}\right\|}^{2}+g(x),$
@@
Meanwhile, since $(y^{k+1},z^{k+1})\in T_{\Lambda}(y^{k},z^{k})$, we have $R(y^{k},z^{k})^{2}\leq{\left\|{(y^{k},z^{k})-(y^{k+1},z^{k+1})}\right\|}^{2}$. These together with (3.24) and (3.55) give the desired result for the case $\beta<L_{{}_{\rm{p}}}^{-1}$. For the other case, we take the summation of (3.24) from 1 to $N$ to obtain
@@ 3.4 Comments on the weak convexity assumption on $g$
$P_{\Lambda}(y,z)=\bigcup_{\begin{subarray}{c}x\in{\rm prox}_{\alpha f}(z)\\ \xi\in\partial p(y)\end{subarray}}\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}\,\Phi_{\Lambda,\xi}(w;y,z,x).$ (3.63)
in Lemma 3.2, but we can actually still ensure Eq. 3.63 without this weak convexity assumption on $g$ in some special cases, such as when either
  1. (I)
    $f=h=0$, or
  2. (II)
    $L_{{}_{\rm{p}}}=0$.
In the former case, ${\rm prox}_{\alpha f}(z)\equiv z$ and $\gamma=\beta$, and consequently $P_{\Lambda}={\rm prox}_{\gamma g}(T_{\gamma p}(y))$, which is precisely the right-hand side of Eq. 3.63. On the other hand, when $L_{{}_{\rm{p}}}=0$, we have $\gamma=\alpha$ and hence
@@
which is clearly equal to the right-hand side of (3.63). We also note that when $g$ is not weakly convex, we can get a weaker version of Lemma 3.3. Indeed, from the definition of $V_{\Lambda,\xi}$ in Eq. 3.7, we can obtain by setting $w=y$ that
@@
    There is $\omega\in(0,\infty)$ (dependent on the stepsizes and the problem) such that the residual converges to 0 with rate
@@
On the other hand, from Eq. 3.3, we get $y^{k+1}-y^{k}=x^{k}-x^{k-1}+\tau^{-1}(z^{k+1}-z^{k})-\tau^{-1}(z^{k}-z^{k-1})$. It follows from the triangle inequality and the estimate $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$ (for $a,b,c\in{\rm I}\!{\rm R}$) that
@@
(3.66)${\leq}\left(3+3\tau^{-2}(1+\alpha L_{{}_{\rm{f}}})^{2}\right)\left(-\frac{2\tau\alpha(V_{0}-\Psi^{*})}{c(\alpha)}\right)+3\tau^{-2}(1+\alpha L_{{}_{\rm{f}}})^{2}\left(-\frac{2\tau\alpha(V_{0}-\Psi^{*})}{c(\alpha)}\right)$ )
$=\left(3+6\tau^{-2}(1+\alpha L_{{}_{\rm{f}}})^{2}\right)\left(-\frac{2\tau\alpha(V_{0}-\Psi^{*})}{c(\alpha)}\right)$
@@
Generalizing the above discussion, provided that the minimization of $\Phi_{\Lambda,\xi^{k}}(w;y^{k},z^{k},x^{k})$ is not difficult or expensive to compute, we can simply replace the $y$-update rule (3.2) in Algorithm 1 with the formula
$y^{k+1}\in\bigcup_{\begin{subarray}{c}x^{k}\in{\rm prox}_{\alpha f}(z^{k})\\ \xi^{k}\in\partial p(y^{k})\end{subarray}}\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}\,\Phi_{\Lambda,\xi^{k}}(w;y^{k},z^{k},x^{k}).$ (3.67)
In this case, we can also apply the results to situations outside conditions (I) and (II) above. Similarly, we can obtain identical estimates for the stepsizes and derive the analogous results as in Proposition 3.19, with $T_{\Lambda}$ in (iii) and in the definition of residual in (3.59) replaced with the set-valued operator on the right-hand side of (3.67).
While finalizing this paper, we became aware of a recent preprint by Dao et al. (2024) that proposed to solve Eq. 1.1 with $L_{{}_{\rm{p}}}=0$ using an algorithm similar to ours. [^1] Specifically, setting $\theta=1$ in (Dao et al., 2024, Algorithm 1) coincides with our algorithm for the case $L_{{}_{\rm{p}}}=0$. Subsequential convergence was proved in (Dao et al., 2024) without the weak convexity assumption on $g$, which is also covered by our framework as it corresponds to case (II) above. We highlight that another advantage of our analysis for this special case in which the settings coincide is that the upper bounds for stepsizes we computed are significantly larger than those in (Dao et al., 2024); see Figs. 2 and 3 for geometric illustrations.
@@ 4 Numerical experiments
In this section, we present experiments to corroborate the theoretical findings in the previous section. All experiments were conducted using MATLAB running on a 64-bit machine with Intel Xeon Silver 4208 CPU and 128GB memory. For our method, we experiment with various values of $\tau$. For each $\tau$, we use the corresponding $\bar{\alpha}$ in Theorems 3.5, 3.9 and 3.12 times $9/10$ to ensure global subsequential convergence as suggested by Theorem 3.18. For comparing different methods that can fit in the framework of Algorithm 1, we report their required number of iterations and running time to reach $R(y,z)\leq 10^{-6}$.

### v2

@@ 3.3 Subsequential convergence
After obtaining the stepsize upper bounds, our next goal is to show subsequential convergence and convergence rates of Algorithm 1. We first establish the boundedness of its iterate sequence.
@@
    $\tau\geq 2$ satisfies Eqs. 3.48 and 3.49, and $\alpha\in\left(\frac{\tau\mu_{*}}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})},\frac{\tau\mu^{*}}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}\right)\cap\left(0,\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right)$.
@@
By our choice of $(\alpha,\beta,\gamma)$, it follows from Theorems 3.5, 3.9 and 3.12 that $\{V_{k}\}$ is strictly decreasing. Meanwhile, we have from Remark 3.6 and Remark 3.11 that under conditions (a) and (b), respectively, it holds that $\alpha<\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$. In condition (c), it is explicitly assumed that $\alpha<\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$. Hence, the second term in (3.55) is always bounded below by zero. By our choice of $\beta$, the last term is also nonnegative. With these, we conclude that $\{\Psi(y^{k})\}$ is bounded above, and therefore $\{y^{k}\}$ is a bounded sequence by the level-boundedness of $\Psi$. On the other hand, since $\Psi$ is closed, $\Psi$ is also bounded below. It thus follows from Eq. 3.55 that $\{{\left\|{y^{k}-x^{k-1}}\right\|}\}$ is bounded above. Since $\{y^{k}\}$ is bounded, it then follows that $\{x^{k-1}\}$ is also bounded. Finally, since $z^{k}=x^{k}+\alpha\nabla f(x^{k})$ by (3.1) and $\nabla f$ is continuous, we obtain the boundedness of $\{z^{k}\}$. This completes the proof of part (i).
@@
$0\in\nabla f(w^{*})+\partial g(w^{*})+\nabla h(w^{*})+\partial p(w^{*}).$ (3.57)
@@
$T_{\Lambda}(y,z)\coloneqq\left\{\begin{pmatrix}y^{\scriptscriptstyle+}\\ z+\tau(y^{\scriptscriptstyle+}-{\rm prox}_{\alpha f}(z))\end{pmatrix}:y^{\scriptscriptstyle+}\in P_{\Lambda}(y,z)\right\}.$ (3.58)
@@
If $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$, then $y^{*}\in P_{\Lambda}(y^{*},z^{*})$, $z^{*}\in z^{*}+\tau(y^{*}-{\rm prox}_{\alpha f}(z^{*}))$. Hence, by noting that ${\rm prox}_{\alpha f}$ is single-valued since $\alpha<L_{{}_{\rm{f}}}^{-1}$, we have
@@
Adding (3.60) and (3.61) then leads to Eq. 3.57. ∎
@@
$\omega\coloneqq\begin{cases}\frac{2\beta(V_{0}-\Psi^{*})}{\min\left\{2\beta\zeta,1-\beta L_{{}_{\rm{p}}}\right\}}&\text{if}~{}\beta<L_{{}_{\rm{p}}}^{-1},\\ \frac{4\gamma(V_{0}-\Psi^{*})}{\min\left\{2\gamma\zeta,1-\gamma\rho_{{}_{\rm{g}}}\right\}}&\text{if}~{}\gamma<\rho_{{}_{\rm{g}}}^{-1},\end{cases}\quad\text{and}\quad\zeta\coloneqq\frac{-c(\alpha)}{2\tau\alpha(1+\alpha L_{{}_{\rm{f}}})^{2}}>0,$
@@
Let $\{(x^{k_{j}},y^{k_{j}},z^{k_{j}})\}$ be such that $(x^{k_{j}},y^{k_{j}},z^{k_{j}})\to(x^{*},y^{*},z^{*})$. From the proof of Proposition 3.15, we have that ${\left\|{y^{k}-x^{k-1}}\right\|}\to 0$ and ${\left\|{y^{k}-y^{k-1}}\right\|}\to 0$. It then follows that $y^{k_{j}}\to x^{*}$ by the triangle inequality, and therefore $x^{*}=y^{*}$. To show that $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$, we first note that since $x^{*}=y^{*}$ and $z^{k}=x^{k}+\alpha\nabla f(x^{k})$ by (2.4), it follows that $z^{*}=y^{*}+\alpha\nabla f(y^{*})$. Since $\alpha<\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\leq\frac{1}{L_{{}_{\rm{f}}}}$, then $f+\frac{1}{2\alpha}\|\cdot\|^{2}$ is strongly convex by Remark 2.2 and therefore $y^{*}\in{\rm prox}_{\alpha f}(z^{*})$. Thus, $z^{*}=z^{*}+\tau(y^{*}-{\rm prox}_{\alpha f}(z^{*}))$. Hence, to show that $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$, it remains to prove that $y^{*}\in P_{\Lambda}(y^{*},z^{*})$. To this end, note that by the convexity and continuity of $\frac{L_{{}_{\rm{p}}}}{2}\|\cdot\|^{2}-p$ on $\operatorname*{dom}(g)$ (by Assumption 1.1(d)) and the formula (3.21), we have from (Beck, 2017, Theorem 3.16) and (3.21) that $\{\xi^{k_{j}}\}$ is bounded. Using again the continuity of $p$ on $\operatorname*{dom}(g)$ together with (2.2) and the fact that $\{y^{k}\}\subseteq\operatorname*{dom}(g)$, we see that accumulation points of $\{\xi^{k_{j}}\}$ belong to $\partial p(y^{*})$. Without loss of generality, we may assume that $\xi^{k_{j}}\to\xi^{*}$ where $\xi^{*}\in\partial p(y^{*})$.
@@
Note that $y^{k_{j}+1}\to y^{*}$ and $u^{k_{j}}\to u^{*}\coloneqq\frac{\gamma}{\alpha}(2y^{*}-z^{*}-\alpha\nabla h(y^{*}))+\frac{\gamma}{\beta}(y^{*}-\beta\xi^{*})$, where we have used the fact that $\xi^{k_{j}}\to\xi^{*}$ and $x^{*}=y^{*}$. To prove the claim that $y^{*}\in P_{\Lambda}(y^{*},z^{*})$, we only need to show that $y^{*}\in{\rm prox}_{\gamma g}(u^{*})$, noting that $y^{*}\in{\rm prox}_{\alpha f}(z^{*})$
@@
$\leq(f+h+p)(\bar{x})+\left\langle\nabla(f+h)(\bar{x})+\bar{\xi},x-\bar{x}\right\rangle+\frac{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+L_{{}_{\rm{p}}}}{2}{\left\|{x-\bar{x}}\right\|}^{2}+g(x),$
@@
Meanwhile, we have $R(y^{k},z^{k})^{2}\leq{\left\|{(y^{k},z^{k})-(y^{k+1},z^{k+1})}\right\|}^{2}$ since $(y^{k+1},z^{k+1})\in T_{\Lambda}(y^{k},z^{k})$. These together with (3.24) and (3.56) give the desired result for the case $\beta<L_{{}_{\rm{p}}}^{-1}$. For the other case, the summation of (3.24) from 1 to $N$ gives
@@ 3.4 Comments on the weak convexity assumption on $g$
$P_{\Lambda}(y,z)=\bigcup_{x\in{\rm prox}_{\alpha f}(z),\xi\in\partial p(y)}\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}\,\Phi_{\Lambda,\xi}(w;y,z,x).$ (3.66)
in Lemma 3.2, but we can actually still ensure Eq. 3.66 without this weak convexity assumption on $g$ in some special cases, such as when either (I) $f=h=0$, or (II) $L_{{}_{\rm{p}}}=0$. In the former case, ${\rm prox}_{\alpha f}(z)\equiv z$ and $\gamma=\beta$, and consequently $P_{\Lambda}={\rm prox}_{\gamma g}(T_{\gamma p}(y))$, which is precisely the right-hand side of Eq. 3.66. On the other hand, when $L_{{}_{\rm{p}}}=0$, we have $\gamma=\alpha$ and hence
@@
which is clearly equal to the right-hand side of (3.66).
We also note that when $g$ is not weakly convex, we can get a weaker version of Lemma 3.3. Indeed, from Eq. 3.7, we can obtain by setting $w=y$ that
@@
    There exists $\omega\in(0,\infty)$ (dependent on the stepsizes and the problem) such that the residual converges to 0 with rate
@@
On the other hand, from Eq. 3.3, we get $y^{k+1}-y^{k}=x^{k}-x^{k-1}+\tau^{-1}(z^{k+1}-z^{k})-\tau^{-1}(z^{k}-z^{k-1})$ It follows from the triangle inequality and the estimate $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$ (for $a,b,c\in{\rm I}\!{\rm R}$) that
@@
(3.69)${\leq}\left(3+6\tau^{-2}(1+\alpha L_{{}_{\rm{f}}})^{2}\right)\left(-\frac{2\tau\alpha(V_{0}-\Psi^{*})}{c(\alpha)}\right).$ ) .
@@
Generalizing the above discussion, provided that the minimization of the function $\Phi_{\Lambda,\xi^{k}}(\cdot,y^{k},z^{k},x^{k})$ is not difficult or expensive to compute, we can simply replace the $y$-update rule (3.2) in Algorithm 1 with the formula
$y^{k+1}\in\bigcup_{x^{k}\in{\rm prox}_{\alpha f}(z^{k}),\xi^{k}\in\partial p(y^{k})}\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}\,\Phi_{\Lambda,\xi^{k}}(w;y^{k},z^{k},x^{k}).$ (3.70)
In this case, we can also apply the results to situations outside conditions (I) and (II) above. Similarly, we can obtain identical estimates for the stepsizes and derive results analogous to Proposition 3.19, where $T_{\Lambda}$ is as defined in (3.58) with $P_{\Lambda}$ given by the right-hand side of (3.66).
While finalizing this paper, we became aware of a recent preprint by Dao et al. (2024) that proposed to solve Eq. 1.1 with $L_{{}_{\rm{p}}}=0$ using an algorithm similar to ours. [^1] Specifically, setting $\theta=1$ in (Dao et al., 2024, Algorithm 1) coincides with our algorithm for the case $L_{{}_{\rm{p}}}=0$. Subsequential convergence was proved in (Dao et al., 2024) without the weak convexity assumption on $g$, which is also covered by our framework, in view of Proposition 3.19, as it corresponds to case (II) above. We highlight that another advantage of our analysis for this special case in which the settings coincide is that the upper bounds for stepsizes we computed are significantly larger than those in (Dao et al., 2024); see Figs. 2 and 3 for geometric illustrations. Meanwhile, (Dao et al., 2024) also proved global convergence of the full sequence under a Kurdyka-Łojasiewicz hypothesis, using the techniques by (Liu et al., 2019b), and a similar proof strategy can be employed to prove the same global convergence property for our algorithm.
@@ 4 Numerical experiments
In this section, we present experiments to corroborate the theoretical findings in the previous section. All experiments were conducted in MATLAB. For our method, we experiment with various values of $\tau$. For each $\tau$, we use the corresponding $\bar{\alpha}$ in Theorems 3.5, 3.9 and 3.12 times $9/10$ to ensure global subsequential convergence as suggested by Theorem 3.18. For comparing different methods that can fit in the framework of Algorithm 1, we report their required number of iterations and running time to reach $R(y,z)\leq 10^{-6}$.